OpenStudy (anonymous):

simplify:a^3 b^2/5a^4b^7*2a^6b^/a^9b^3

OpenStudy (anonymous):

you have made mistake 6^/

OpenStudy (anonymous):

its a^6 b

OpenStudy (anonymous):

\[\frac{\frac{a^3b^2}{5a^4b^7*2a^6b}}{a^9b^3}?\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

that looks correct to what he has said?

OpenStudy (anonymous):

\[\frac{a^3b^2}{5a^4b^7*2a^6b}*\frac{1}{a^9b^3}\]

OpenStudy (anonymous):

\[\frac{a^3b^2}{5a^4b^7}\cdot\frac{2a^6b}{a^9b^3}=\frac{1}{5a^4b^7}\]

OpenStudy (anonymous):

that is correct for if it is 2 fractions

OpenStudy (anonymous):

sorry missed 2 (WHERES EDIT BUTTON), answer is \[\frac{2}{5a^4b^7}\]

OpenStudy (anonymous):

he placed only 1 asterisk (not after every a*b*a*b*...) so we can assume it's 2 fractions

OpenStudy (anonymous):

Hey ocuane, next time you do such a problem, it'd be helpful to use the Equation Editor button below the text box.

OpenStudy (anonymous):

yes tomas but it can be written such as what i put above. It doesn't necessarily need * every single time

OpenStudy (anonymous):

he'd need to put it in parenthesis

OpenStudy (anonymous):

\[5*a*b*6*a*b=5ab*6ab\]

OpenStudy (anonymous):

i know that it can be interpreted various ways

myininaya (myininaya):

that is hard to read

myininaya (myininaya):

i think tomas does have the right interpretation

myininaya (myininaya):

but i could be wrong